I have a space that is 6 feet wide by 7 feet long. I need to know if it is enough room to build a 100 square foot greenhouse?
Answer by Alan3354(69443)

I have a space that is 6 feet wide by 7 feet long. I need to know if it is enough room to build a 100 square foot greenhouse?
------------
If it's a rectangle, area = L*W = 6*7
Area = 42 sq ft
Less than 100 sf